# Concept-based Adversarial Attack: a Probabilistic Perspective

### v1

@@ Appendix I Datasets and Licenses
$\uparrow$ HyperIQA 0.7255 0.6486 0.5471 0.6808 0.6998 0.6865
$\uparrow$ DBCNN 0.6956 0.6175 0.5236 0.6310 0.6577 0.6402
$\uparrow$ ARNIQA 0.7667 0.6909 0.6872 0.7112 0.7328 0.7142
$\uparrow$ MUSIQ-AVA 4.3760 4.3042 4.0496 4.3236 4.6125 4.5700
$\uparrow$ NIMA-AVA 4.5595 4.4532 4.0549 4.5317 4.8172 4.7778
$\uparrow$ MUSIQ-KonIQ 65.0549 59.5258 51.9209 59.7813 63.5363 61.6641
$\uparrow$ TReS 93.2127 86.2644 73.0392 86.4376 90.7757 88.9947
Table 9: Targeted attack success rates (in %) on ImageNet classifiers, with ViT-Base serving as the white-box victim (surrogate) classifier.
ACA DiffAttack ProbAttack OURS (CONS) OURS (AGGR)
White-box Top1
ViT-Base 3.46 81.41 63.33 85.51 85.51
Transferability Top100
VGG19 13.46 20.38 12.05 16.92 23.33
ResNet 152 15.13 26.79 15.26 18.59 26.92
DenseNet 161 14.74 28.46 18.59 21.03 33.21
Inception V3 12.18 21.54 13.33 13.97 20.00
EfficientNet B7 16.03 24.62 14.23 19.23 26.28
Adversarial Defence Top100
Inception V3 Adv 15.64 21.41 16.41 16.79 24.36
EfficientNet B7 Adv 13.59 22.18 16.15 20.26 31.79
Ensemble IncRes V2 13.97 18.33 13.46 15.00 22.31
@@
Clean ACA DiffAttack ProbAttack OURS (CONS) OURS (AGGR)
@@
$\uparrow$ Avg. Clip Score 1.0 0.7681 0.8040 0.8586 0.8222 0.8104
@@
$\uparrow$ HyperIQA 0.7255 0.6324 0.5475 0.6796 0.7077 0.6961
$\uparrow$ DBCNN 0.6956 0.5842 0.5222 0.6254 0.6652 0.6526
$\uparrow$ ARNIQA 0.7667 0.6765 0.6844 0.7120 0.7351 0.7195
$\uparrow$ MUSIQ-AVA 4.3760 4.3131 4.0400 4.2450 4.5401 4.5345
$\uparrow$ NIMA-AVA 4.5595 4.4775 4.0321 4.4918 4.7706 4.7747
$\uparrow$ MUSIQ-KonIQ 65.0549 57.6105 52.0604 59.3605 64.2536 62.4264
$\uparrow$ TReS 93.2127 83.6834 73.4362 85.8749 91.7463 89.7815

### v2

@@ Appendix I Compute Resources
$\uparrow$ HyperIQA 0.7255 0.5002 0.6486 0.5471 0.6808 0.6998 0.6865
$\uparrow$ DBCNN 0.6956 0.5040 0.6175 0.5236 0.6310 0.6577 0.6402
$\uparrow$ ARNIQA 0.7667 0.5972 0.6909 0.6872 0.7112 0.7328 0.7142
$\uparrow$ MUSIQ-AVA 4.3760 3.7912 4.3042 4.0496 4.3236 4.6125 4.5700
$\uparrow$ NIMA-AVA 4.5595 3.7879 4.4532 4.0549 4.5317 4.8172 4.7778
$\uparrow$ MUSIQ-KonIQ 65.0549 49.9072 59.5258 51.9209 59.7813 63.5363 61.6641
$\uparrow$ TReS 93.2127 63.7643 86.2644 73.0392 86.4376 90.7757 88.9947
@@
Clean NCF ACA DiffAttack ProbAttack OURS (CONS) OURS (AGGR)
@@
$\uparrow$ Avg. Clip Score 1.0 0.8733 0.7681 0.8040 0.8586 0.8222 0.8104
@@
$\uparrow$ HyperIQA 0.7255 0.5006 0.6324 0.5475 0.6796 0.7077 0.6961
$\uparrow$ DBCNN 0.6956 0.5027 0.5842 0.5222 0.6254 0.6652 0.6526
$\uparrow$ ARNIQA 0.7667 0.5879 0.6765 0.6844 0.7120 0.7351 0.7195
$\uparrow$ MUSIQ-AVA 4.3760 3.7822 4.3131 4.0400 4.2450 4.5401 4.5345
$\uparrow$ NIMA-AVA 4.5595 3.7666 4.4775 4.0321 4.4918 4.7706 4.7747
$\uparrow$ MUSIQ-KonIQ 65.0549 50.0413 57.6105 52.0604 59.3605 64.2536 62.4264
$\uparrow$ TReS 93.2127 64.1109 83.6834 73.4362 85.8749 91.7463 89.7815
Table 18: Quantitative comparison of similarity to the original images and no reference image quality metrics for unrestricted adversarial examples with ConvNeXT serving as the victim (surrogate) classifier.
Clean NCF ACA DiffAttack ProbAttack OURS (CONS) OURS (AGGR)
Similarity
$\uparrow$ Avg. Clip Score 1.0 0.8621 0.7542 0.7964 0.8467 0.8139 0.8011
Image Quality
$\uparrow$ HyperIQA 0.7255 0.4928 0.6281 0.5392 0.6703 0.6933 0.6810
$\uparrow$ DBCNN 0.6956 0.4951 0.5784 0.5179 0.6201 0.6531 0.6405
$\uparrow$ ARNIQA 0.7667 0.5835 0.6710 0.6801 0.7066 0.7294 0.7118
$\uparrow$ MUSIQ-AVA 4.3760 3.7511 4.2814 4.0206 4.2314 4.5116 4.5030
$\uparrow$ NIMA-AVA 4.5595 3.7524 4.4433 4.0215 4.4830 4.7464 4.7492
$\uparrow$ MUSIQ-KonIQ 65.0549 49.3812 57.0114 51.7324 58.9012 63.1025 61.3571
$\uparrow$ TReS 93.2127 62.9011 82.7410 72.1109 85.2442 90.9211 88.7724
Table 19: Quantitative comparison of similarity to the original images and no reference image quality metrics for unrestricted adversarial examples with ResNet-50 Adv. serving as the victim (surrogate) classifier.
Clean NCF ACA DiffAttack ProbAttack OURS (CONS) OURS (AGGR)
Similarity
$\uparrow$ Avg. Clip Score 1.0 0.8556 0.7485 0.7877 0.8367 0.8059 0.7945
Image Quality
$\uparrow$ HyperIQA 0.7255 0.4854 0.6221 0.5297 0.6617 0.6875 0.6750
$\uparrow$ DBCNN 0.6956 0.4908 0.5718 0.5070 0.6078 0.6473 0.6342
$\uparrow$ ARNIQA 0.7667 0.5764 0.6654 0.6683 0.6932 0.7231 0.7070
$\uparrow$ MUSIQ-AVA 4.3760 3.7435 4.2637 4.0130 4.2223 4.5031 4.4966
$\uparrow$ NIMA-AVA 4.5595 3.7453 4.4342 4.0113 4.4676 4.7399 4.7351
$\uparrow$ MUSIQ-KonIQ 65.0549 49.3740 57.0028 51.7202 58.1513 63.0946 61.3527
$\uparrow$ TReS 93.2127 62.8961 82.7340 72.1034 84.3081 90.4438 88.0786